OpenStudy (anonymous):

suppose (-8,0) and (8,0) are the foci of an ellipse. using the distance formula write an equation stating that the sum of the two distances from a point p(x,y) on the ellipse to the two foci is 20

OpenStudy (mertsj):

\[\sqrt{(x+8)^2+(y-0)^2}+\sqrt{(x-8)^2+(y-0)^2}=20\]

OpenStudy (anonymous):

now eliminate the radicals in this equation by squaring twice. simplify the resulting equation and write it in standard form of the equation of an ellipse. what are a and b for this ellipse?

OpenStudy (mertsj):

Did you do it?

OpenStudy (mertsj):

\[\sqrt{(x+8)^2+y^2}=20-\sqrt{(x-8)^2+y^2}\] \[(x+8)^2+y^2=400-40\sqrt{(x-8)^2+y^2}+(x-8)^2+y^2\] \[x^2+16x+64+y^2=400-40\sqrt{(x-8)^2+y^2}+x^2-16x+64+y^2\] \[32x-400=-40\sqrt{(x-8)^2+y^2}\] \[4x-50=-5\sqrt{(x-8)^2=y^2}\] \[16x^2-400x+2500=25[(x-8)^2+y^2]\] \[16x^2-400x+2500=25(x^2-16x+64+y^2)\] \[16x^2-400x+2500=25x^2-400x+1600+25y^2\] \[900=9x^2+25y^2\]

OpenStudy (mertsj):

Divide both sides by 900

OpenStudy (mertsj):

\[\frac{x^2}{100}+\frac{y^2}{36}=1\]

OpenStudy (mertsj):

a=10, b=6